a man is 37 years old and his child's age is 8. how many year ago was the product of their ages 96
Answer by stanbon(75887)

a man is 37 years old and his child's age is 8. how many year ago was the product of their ages 96
---------------------
Solve: (37-x)(8-x) = 96
------------
296 - 35x + x^2 = 96
------
x^2 -45x + 200 = 0
------
x = 5 (# of yrs. ago when man is 32 and child is 3)
=================================
Cheers,
Stan H.